OpenStudy (anonymous):

Jerry's bowling score was 230 last week, but only 190 this week. What was the percent decrease in Jerry's score? A. 21.1% B. 17.4% C. 1.74% D. 2.11%

OpenStudy (anonymous):

@Thesmarterone

OpenStudy (anonymous):

@Thesmarterone need help in math me think it D

OpenStudy (anonymous):

230*0.174=40.02 230-40.02=189.98

OpenStudy (anonymous):

You have to devide the new score by the old one: 190/230= 0,826 Now subtract 1 0,826-1= -0,173 Take the module and multiply by 100 17,39% Latter b

OpenStudy (anonymous):

yes, 17.4 is the closest answer :)

OpenStudy (anonymous):

I was never good with percents I don't know what happened to me now I'm a teenager I think I don't need school and it sucks